24589, 197084, 961582, ? find the?
Answer by Edwin McCravy(20056)

24589, 197084, 961582, ? find the?
a(n) = 296001.5n²-715509.5n+444097

Proof:

a(1) = 296001.5(1)²-715509.5(1)+444097
a(1) = 296001.5-715509.5+444097
a(1) = 24589

a(2) = 296001.5(2)²-715509.5(2)+444097
a(2) = 296001.5(4)-1431019+444097
a(2) = 1184006-1431019+444097
a(2) = 197084

a(3) = 296001.5(3)²-715509.5(3)+444097
a(3) = 296001.5(9)-2146528.5+444097
a(3) = 2664013.5-2146528.5+444097
a(3) = 961582

a(4) = 296001.5(4)²-715509.5(4)+444097
a(4) = 296001.5(16)-2862038+444097
a(4) = 4736024-2862038+444097
a(4) = 2318083

Edwin